40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 29 19:50:49 2013, Jenne, Update, Locking, Better POP layout, no extra PRM motion with locked cavity HalfPRCL_PRM-flatMirror_RefsAreLocked_OthersUnlocked.pdfHalfPRCL_PRM-flatMirror_RefsAreLocked_OthersUnlocked.png
    Reply  Tue Jan 29 21:07:48 2013, Jenne, Update, Locking, PRM coils need diagonalizing 
       Reply  Wed Jan 30 19:34:17 2013, Jenne, Update, Locking, PRM coils need diagonalizing 
    Reply  Wed Jan 30 19:30:21 2013, Jenne, Update, Locking, POP in-vac improved HalfPRCL_PRM-flatMirror_POPtrans_30Jan2013.jpg
Message ID: 7957     Entry time: Tue Jan 29 19:50:49 2013     Reply to this: 7959   7968
Author: Jenne 
Type: Update 
Category: Locking 
Subject: Better POP layout, no extra PRM motion with locked cavity 

[Jenne, Jamie, Manasa]

Today's activities focused on getting the POP layout improved, so that we could get clean data for the mode scan measurement. 

As Jamie and Koji pointed out yesterday, the beam was still a little too big on the POP DC PD, and was falling off the diode when the beam moved a small amount.  We have fixed things so that the PD is now at the focus of the lens, and the camera is at a place where the beam takes up most of the area on the TVs.  The beam no longer falls off the PD with cavity fluctuations.  A key point of this work was also to use an extra 2" optic to steer the beam down the length of the POP table, and then do the 50/50 beam splitting later with a 1" optic.  The 1" BS that we had been using (including with the "real" POP beam) is too small.  We could not find a 2" 50/50 BS, so we opted to do the splitting closer to the focal point.  Also, the BS that was splitting the beam between the PD and the camera was a 33% reflector, but now is a 50/50 BS. When we put back the 'real' POP path, we need to consider using larger optics, or a faster lens. The POP path is now good, hopefully for the duration of the half cavity test.

After getting the POP path taken care of, and tweaking up the cavity alignment a little bit, the transmitted power on POP DC is ~22,000 counts, with occasional fluctuations as high as 25,000 counts. 

Jamie looked at the REFL path, and things look sensible there.  The unlocked REFL power is ~36 counts, and the locked power is ~20 counts.  I'm not sure what the 160 counts that Koji mentioned in his edits to elog 7949 is about.

I looked at the PRM oplev with the cavity locked and unlocked, and with today's alignment, there seems to be no difference in the amount of PRM motion when the cavity is locked vs unlocked. 

HalfPRCL_PRM-flatMirror_RefsAreLocked_OthersUnlocked.png

 It still looks like we might be seeing some clipping in the in-vac POP steering mirrors - we haven't gotten to them yet.

Jamie is currently modifying Yuta's mode scan analysis script to look at the data that we have of the cavity.

 


We need more 2" optics.  There are no mounted 2" spares in the various optic "graveyards" (which, PS, we should consolidate all into the cabinet with doors near the optics bench), and the options for boxes in the drawers is slim pickin's.  We have some S-pol stuff, but no Y1s or BS-50s for P-pol.  Since POP, POX, POY, IPANG, TRX and TRY all come out of the vacuum with large beams, we should have some options for these laying around for this kind occasional temporary thing.  We also need to choose, then purchase better 2" lenses for the pickoffs.

Attachment 1: HalfPRCL_PRM-flatMirror_RefsAreLocked_OthersUnlocked.pdf  65 kB  | Hide | Hide all
HalfPRCL_PRM-flatMirror_RefsAreLocked_OthersUnlocked.pdf
ELOG V3.1.3-